# Trajectory Consistency Distillation

### v1

@@ Abstract
Latent Consistency Model (LCM) extends the Consistency Model to the latent space and leverages the guided consistency distillation technique to achieve impressive performance in accelerating text-to-image synthesis. However, we observed that LCM struggles to generate images with both clarity and detailed intricacy. To address this limitation, we initially delve into and elucidate the underlying causes. Our investigation identifies that the primary issue stems from errors in three distinct areas. Consequently, we introduce Trajectory Consistency Distillation (TCD), which encompasses trajectory consistency function and strategic stochastic sampling. The trajectory consistency function diminishes the distillation errors by broadening the scope of the self-consistency boundary condition and endowing the TCD with the ability to accurately trace the entire trajectory of the Probability Flow ODE. Additionally, strategic stochastic sampling is specifically designed to circumvent the accumulated errors inherent in multi-step consistency sampling, which is meticulously tailored to complement the TCD model. Experiments demonstrate that TCD not only significantly enhances image quality at low NFEs but also yields more detailed results compared to the teacher model at high NFEs.
@@ 1 Introduction
To overcome this limitation, Song et al. (2023) proposed Consistency Models (CMs), an emerging class of powerful generative models capable of generating high-quality data with single-step or few-step sampling without the need for adversarial training. CMs are inextricably connected with SGMs in their underlying mathematical underpinnings, aiming to enforce the self-consistency property by mapping arbitrary points on the trajectory of the same Probability Flow Ordinary Differential Equation (PF ODE) to the trajectory’s origin (Song et al., 2020b). CMs can be trained with consistency distillation or treated as standalone generative models. Song et al. (2023) have demonstrated their superiority through extensive experiments in the pixel space. Latent Consistency Models (LCMs) (Luo et al., 2023a) further incorporate Latent Diffusion Models (LDMs) (Rombach et al., 2022) with CMs, achieving remarkable success in swiftly synthesizing high-resolution images conditioned on text. Moreover, LCM-LoRA (Luo et al., 2023b) improves the training efficiency of LCMs and converts them into a universal neural PF ODE solver by introducing LoRA (Hu et al., 2021) into the distillation process of LCMs. It is noteworthy that all these Consistency-Type Models still allow for striking a balance between computation and sample quality using Multistep Consistency Sampling (Song et al., 2023). In particular, allocating additional compute for more iterations can theoretically yield samples of higher quality.
@@
Despite the introduction of the guided distillation method and skipping-step technique (Luo et al., 2023a) by LCMs for effectively distilling knowledge from pre-trained diffusion models, the quality of images generated by LCMs in a single step or even with minimal steps (4$\sim$8) still lags significantly behind the convergence of its teacher model. Our investigation revealed that, in practice, increasing the number of inference iterations diminishes the visual complexity and quality of the results, as illustrated in Figure 1. This renders LCMs less capable of synthesizing samples that are perceptually comparable to those of their teacher model. Recent findings by EDM (Karras et al., 2022) identified that an excessive Langevin-like addition and removal of random noise result in a gradual loss of detail in the generated images. Additionally, Li et al. (2023) showed evidence that during the process of multi-step sampling, discrete errors accumulate over iterations, ultimately causing the generated images to deviate from the target distribution.
Drawing inspiration from the observation, we first meticulously examine the training procedure and delve into the multistep consistency sampling procedure to figure out the root cause. Our investigation revealed that the issue stems from the cumulative errors inherent in multi-step sampling processes. These errors predominantly originate from three sources: 1) the estimation errors in the original score matching model, 2) the distillation errors in the consistency model, and 3) the discretisation errors accumulated during the sampling phase. Collectively, these errors significantly undermine the efficacy of the multistep sampling consistency model, resulting in a performance that falls substantially short of expectations.
To suppress these errors, we introduce Trajectory Consistency Distillation (TCD), as summarized in Figure 2, which comprises two key elements: trajectory consistency function and strategic stochastic sampling. Specifically, inspired by the form of exponential integrators, the trajectory consistency function (TCF) diminishes distillation errors by expanding the boundary conditions of the consistency model and enabling seamless transitions at any point along the trajectory governed by the PF ODE, as demonstrated in Figure 1(a). Furthermore, strategic stochastic sampling (SSS) suppresses the accumulated discretisation error and estimation error according to the narrowed bi-directional iterations as shown in Figure 1(b).
@@ 2.1 Diffusion Models
(b) Sampling process, as detailed in Algorithm 3 & Algorithm 4, where SSS significantly reduces accumulated error through the iterative traversal with the stochastic parameter $\gamma$ compared to the multistep consistency sampling.
@@ 2.2 Consistency Models
$\bm{f}_{\bm{\theta}}(\tilde{\bm{x}}_{t},t)=\begin{cases}\bm{x}_{0},&t=0\\ \texttt{Solver}(F_{\bm{\theta}}(\tilde{\bm{x}}_{t},t),t,0;\bm{\theta}),&t\in(0,T]\end{cases}$
The $\texttt{Solver}(\cdot,t,0;\theta)$ is the update function of an ODE solver used to estimate ${\bm{x}}_{0}$ when given the output $F_{\theta}(\tilde{\bm{x}}_{t},t)$ from a pretrained model at timestep $t$. For training consistency models, the objective of Consistency Distillation (CD) is defined as minimizing:
@@
where $0=t_{1}<t_{2}\cdots<t_{N}=T$, $n$ uniformly distributed over $\{1,2,\cdots,N-1\}$, $\lambda(\cdot)\in\mathbb{R}^{+}$ is a positive weighting function, and the expectation is taken with respect to $\bm{x}\sim p_{\rm{data}}$. $\bm{x}_{t_{n+1}}$ can be sampled using SDE  1 and $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}$ is calculated by $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}:=\Phi(\bm{x}_{t_{n+1}},t_{n+1},t_{n};\bm{\phi})$, where $\Phi(\cdots;\bm{\phi})$ represents the update function of a one-step ODE solver applied to the empirical PF ODE 3. Moreover, $\bm{\theta^{-}}$ is introduced to stabilize the training process and updated by an exponential moving average (EMA) strategy, i.e., given $10\leq\mu<$ 1, $\bm{\theta}^{-}\leftarrow\texttt{sg}(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta})$.
@@
3 Elucidating Errors in Consistency Models
In this section, we elucidate various types of errors that occurred in multistep sampling of consistency models, which motivates us to propose corresponding solutions aimed at refining them. Our analysis reveals that the error is primarily composed of three components: distillation error of consistency models, estimation error of score matching models, and accumulated discretisation error (also known as truncation error) during solving the ODEs.
3.1 Consistency Distillation Error
With a well-trained model $\bm{\theta^{*}}$ such that $\mathcal{L}_{\rm{CD}}^{N}(\bm{\theta^{*}},\bm{\theta^{*}};\bm{\phi})=0$, and considering that the ODE solver $\Phi(\cdots;\bm{\bm{\phi}})$ has the local discretisation error, Theorem 1 in (Song et al., 2023) shows that the consistency distillation error is bounded:
$\mathop{\textnormal{sup}}\limits_{n,{\bm{x}}}\|{\bm{f}}_{\bm{\theta^{*}}}({\bm{x}},t_{n}),{\bm{f}}({\bm{x}},t_{n};\bm{\phi})\|_{2}=\mathcal{O}\left((\Delta t)^{p}\right),$ (7)
with $\Delta t$ and $p$ as defined in (Song et al., 2023).
3.2 Error Bound in Multistep Consistency Sampling
Theoretically, with a well-trained consistency model ${\bm{f}}_{\bm{\theta^{*}}}$, one can generate samples with just one forward pass through the consistency model. However, the one-step sampling yields suboptimal results (Luo et al., 2023a; Song et al., 2023). Thus, multistep consistency sampling has been introduced in (Song et al., 2023) to enhance sample quality through alternating denoising and noise injection steps. We term it as multistep sampling in this manuscript.
For simplicity, we consider VP SDEs in this paper; thus, the drift and diffusion coefficients $\mu(t)$ and $\nu(t)$ in SDE( Equation 1) can be written as:
$\mu(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t},\quad\nu(t)=\sqrt{\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}},$ (8)
where $\alpha_{t}$ and $\sigma_{t}$ specify the noise schedule in the perturbation kernels,
$q_{0t}({\bm{x}}_{t}|{\bm{x}}_{0})=\mathcal{N}({\bm{x}}_{t}|\alpha_{t}{\bm{x}}_{0},\sigma_{t}^{2}\bm{I}).$ (9)
Given a sequence of $N$ sampling timesteps $T=\tau_{1}>\tau_{2}>\cdots>\tau_{N}$ and an initial value $\tilde{{\bm{x}}}_{\tau_{1}}\sim\mathcal{N}(0,\bm{I})$, the generating procedure of the $n$-th step can be written as:
${\bm{x}}_{\tau_{1}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{1}},T),$ (10)
Diffuse: $\tilde{{\bm{x}}}_{\tau_{n}}\leftarrow\alpha_{\tau_{n}}{\bm{x}}_{\tau_{(n-1)}\rightarrow 0}+\sigma_{\tau_{n}}{\bm{z}},\quad{\bm{z}}\sim\mathcal{N}(0,\bm{I})$
Denoise: ${\bm{x}}_{\tau_{n}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{n}},\tau_{n}),$
this process is detailed in Algorithm 3. We derive a further corollary from Corollary 7 in  (Lyu et al., 2023) as follows:
Corollary 3.1.
With the sampling process defined in Equation 10 and denoting the distribution of ${\bm{x}}_{\tau_{n}\rightarrow 0}$ as $q_{\bm{\theta^{*}},n}$, we have single-step sampling result, $q_{\bm{\theta^{*}},1}={\bm{f}}_{{\bm{\theta^{*}}},T}\sharp\mathcal{N}(0,\bm{I})$ and multistep sampling result, $q_{\bm{\theta^{*}},N}={\bm{f}}_{{\bm{\theta^{*}}},\tau_{N}}\sharp\mathcal{N}(\alpha_{\tau_{N}}{\bm{x}}_{\tau_{(N-1)}\rightarrow 0},\sigma_{\tau_{N}}\bm{I})$ with $N>1$. The Total Variational (TV) distance between $q$ and $p_{\rm{data}}$ is
$TV(q_{\bm{\theta^{*}},1},p_{\rm{data}})$ $=\mathcal{O}\left(T(\varepsilon_{cd}+\mathcal{L}_{f}\varepsilon_{se})\right)$
$TV(q_{\bm{\theta^{*}},N},p_{\rm{data}})$ $=\mathcal{O}\left({2^{-N}T(\varepsilon_{cd}+\mathcal{L}_{f}\varepsilon_{se})}\right),$
where normal-$\sharp$ is a push-forward operator associated with a measurable map, $\mathcal{L}_{f}$ is the Lipschitz constant of consistency model, and $\varepsilon_{cd}$, $\varepsilon_{se}$ represent the error of consistency distillation in Equation 7 and score matching estimation, respectively.
The detailed proof can be found in Appendix D.1. From  Corollary 3.1, we observe that multistep sampling has a lower error bound than a single step. This observation aligns with the empirical findings presented in (Karras et al., 2022): the local truncation error scales superlinearly with respect to step size, thus increasing $N$ improves the accuracy of the solution. Moreover, it illustrates that the model performance tends to underperform single-step generation methods (Sauer et al., 2023; Yin et al., 2023) without the help of additional supervision.
3.3 Accumulated Error in Multistep Sampling
However, due to the error of estimation and discretisation in practice, the total error will be accumulated in every sampling step. Denote the distribution of ${\bm{x}}_{\tau_{n}\rightarrow 0}$ output by the true consistency function ${\bm{f}}(\cdot,\cdot;\bm{\phi})$ as $p_{n}$, Theorem 2 in (Chen et al., 2022) demonstrates that the TV error between $q_{\bm{\theta^{*}},n}$ and $p_{n}$ is
$TV(q_{\bm{\theta^{*}},n},p_{n})=\mathcal{O}(\sqrt{\tau_{n}}),$ (11)
the error in each $n$-th step will accumulate over $N$ sampling steps, resulting in the defined accumulated error
$TV(q_{\bm{\theta^{*}},N},p_{N})=\mathcal{O}\left(\sum_{n=1}^{N}\sqrt{\tau_{n}}\right).$ (12)
The above result is a special case of Theorem 4.2, and its proof is presented in Section D.3. As a consequence, there is a significant reduction in image details at higher NFEs, as shown in Figure 3.

### v2

@@ Abstract
Latent Consistency Model (LCM) extends the Consistency Model to the latent space and leverages the guided consistency distillation technique to achieve impressive performance in accelerating text-to-image synthesis. However, we observed that LCM struggles to generate images with both clarity and detailed intricacy. Consequently, we introduce Trajectory Consistency Distillation (TCD), which encompasses trajectory consistency function and strategic stochastic sampling. The trajectory consistency function diminishes the parameterisation and distillation errors by broadening the scope of the self-consistency boundary condition with trajectory mapping and endowing the TCD with the ability to accurately trace the entire trajectory of the Probability Flow ODE in semi-linear form with an Exponential Integrator. Additionally, strategic stochastic sampling provides explicit control of stochastic and circumvents the accumulated errors inherent in multi-step consistency sampling. Experiments demonstrate that TCD not only significantly enhances image quality at low NFEs but also yields more detailed results compared to the teacher model at high NFEs.
@@ 1 Introduction
To overcome this limitation, Song et al. (2023) proposed Consistency Models (CMs), an emerging class of powerful generative models capable of generating high-quality data with single-step or few-step sampling without the need for adversarial training. CMs are inextricably connected with SGMs in their underlying mathematical underpinnings, aiming to enforce the self-consistency property by mapping arbitrary points on the trajectory of the same Probability Flow Ordinary Differential Equation (PF ODE) to the trajectory’s origin (Song et al., 2020b). CMs can be trained with consistency distillation or treated as standalone generative models. Song et al. (2023) have demonstrated their superiority through extensive experiments in the pixel space. Latent Consistency Models (LCMs) (Luo et al., 2023a) further incorporate Latent Diffusion Models (LDMs) (Rombach et al., 2022) with CMs, achieving remarkable success in swiftly synthesizing high-resolution images conditioned on text. Moreover, LCM-LoRA (Luo et al., 2023b) improves the training efficiency of LCMs and converts them into a universal neural PF ODE solver by introducing LoRA (Hu et al., 2021) into the distillation process of LCMs. It is noteworthy that all these Consistency-Type Models still allow for striking a balance between computation and sample quality using Multistep Consistency Sampling (Song et al., 2023). There are also some works that focus on changing the interval of the consistency function or the distillation learning protocol (Gu et al., 2023; Berthelot et al., 2023; Heek et al., 2024), leading to enhanced outcomes.
@@
Despite the introduction of the guided distillation method and skipping-step technique (Luo et al., 2023a) by LCMs for effectively distilling knowledge from pre-trained diffusion models, the quality of images generated by LCMs in a single step or even with minimal steps (4$\sim$8) still lags significantly behind the convergence of its teacher model. Our investigation revealed that, in practice, increasing the number of inference iterations diminishes the visual complexity and quality of the results, as illustrated in Figure 1. This renders LCMs less capable of synthesizing samples that are perceptually comparable to those of their teacher model. Recent findings by EDM (Karras et al., 2022) identified that an excessive Langevin-like addition and removal of random noise result in a gradual loss of detail in the generated images. Additionally, Li et al. (2023) showed evidence that during the process of multi-step sampling, discretisation errors accumulate over iterations, ultimately causing the generated images to deviate from the target distribution. Recently, in order to balance the inference steps and the sample quality during inference, CTM (Kim et al., 2023) proposed a unified framework that takes advantage of the sampling methods of Distillation Models (Salimans & Ho, 2022; Song et al., 2023) and SGMs according to a novel representation of denoisers and analyzed the errors during sampling.
In this paper, we focus on improving the generation quality of LCMs, especially reducing the errors arising during parameterisation process of LCMs. We introduce Trajectory Consistency Distillation (TCD), as summarized in Figure 2, which comprises two key elements: trajectory consistency function and strategic stochastic sampling. Specifically, inspired by the form of exponential integrators (Lu et al., 2022a, b), the trajectory consistency function (TCF) diminishes parameterisation and distillation errors by expanding the boundary conditions of the consistency model and enabling seamless transitions at any point along the trajectory governed by the PF ODE, as demonstrated in Figure 2(a). Furthermore, strategic stochastic sampling (SSS) allows control of stochasticity and suppresses the accumulated discretisation error and estimation error according to the narrowed bi-directional iterations as shown in Figure 2(b).
@@
2 Related Works
Diffusion Models.
Diffusion Models (DMs), also known as Score-based generative models, have shown superior performance in various generative fields. Chen et al. (2022) provides theoretical convergence guarantees, implying that DMs can efficiently sample from essentially any realistic data distribution under minimal data assumptions. ADM (Dhariwal & Nichol, 2021) firstly shows the potential of DMs to outperform GANs. EDM (Karras et al., 2022) further elucidates the design space of DMs, clearly separating the concrete design choices, deriving best practices for the sampling process, and improving the training dynamics, thereby drastically improving the results.
Text-conditional Diffusion Models.
Fast Sampling of DMs.
DMs exhibit great generative capabilities but are bottlenecked by their slow sampling speed. Various numerical methods driven by the SDE mathematical underpinnings underlying DMs have been proposed for accelerating DM sampling. DDIM (Song et al., 2020a) has originally shown promise for few-step sampling. Other works involve predictor-corrector samplers (Song et al., 2020b; Karras et al., 2022), exponential integrators (Lu et al., 2022a; Zhang & Chen, 2022; Lu et al., 2022b), and automated methods to adjust solvers (Kong & Ping, 2021; Jolicoeur-Martineau et al., 2021). Among them, the semi-linear structure utilizing exponential integrators (Lu et al., 2022a; Zhang & Chen, 2022; Lu et al., 2022b) can derive accurate solutions for the PF-ODEs. By separating the linear and nonlinear parts, it skillfully bypasses the inherent discretization errors in the linear items, thereby obtaining exact analytical solutions for the linear term. Another series of research works exemplified by distillation techniques (Luhman & Luhman, 2021; Salimans & Ho, 2022; Meng et al., 2023). These methods distill knowledge from pretrained models into few-step samplers, representing an efficient solution for few NFEs. However, they may experience a lengthy and costly process that requires huge amounts of data and suffer from slow convergence.
Consistency Models.
To overcome the limitations of current fast samplers, Song et al. (2023) proposes Consistency Models (CMs) learning a direct mapping from noise to data built on the top of the trajectory of PF ODE, achieving one-step generation while allowing multi-step sampling to trade compute for quality. Lyu et al. (2023) provide the first convergence guarantees for CMs under moderate assumptions. Kim et al. (2023) proposes a universal framework for CMs and DMs samplings with an “anytime to anytime” decoder, which is able to predict both infinitesimally small steps and arbitrary long step jumps. Together with a $\gamma$-sampler, it achieves a balance between sampling quality and computational resources. Both CTM and TCD eliminate constraints on predictions at the origin during distillation, but from distinct viewpoints, leading to different solutions. Furthermore, CTM provides $\gamma$ sampling with a well-founded theoretical basis in the view of multistep sampling, while our experiments disseminate its theoretical findings. Further comments about CTM and TCD can be found in Appendix A. Latent Consistency Models (LCMs) (Luo et al., 2023a) integrate consistency distillation with latent diffusion models (Rombach et al., 2022) to achieve impressive performance in accelerating text-to-image synthesis. LCM-LoRA (Luo et al., 2023b) further improves the training efficiency and versatility of LCMs by introducing LoRA (Hu et al., 2021) into the distillation process. Several studies have concentrated on modifying the interval of the consistency function and refining the distillation learning protocol (Gu et al., 2023; Berthelot et al., 2023; Heek et al., 2024), resulting in improved results.
Parameter-Efficient Fine-Tuning.
@@ 3.2 Consistency Models
$\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)=\begin{cases}\bm{x}_{0},&t=0\\ F_{\bm{\theta}}(\bm{x}_{t},t),&t\in(0,T]\end{cases}$ (6)
or using skip-connection as:
${\bm{f}}_{\bm{\theta}}({\bm{z}}_{t},{\bm{c}},t)=c_{\text{skip}}(t){\bm{z}}_{t}+c_{\text{out}}(t)F_{\bm{\theta}}(\bm{x}_{t},t),$ (7)
where $c_{\text{skip}}(\epsilon)=1$ and $c_{\text{out}}(\epsilon)=0$.
For training consistency models, the objective of Consistency Distillation (CD) is defined as minimizing:
@@
where $0=t_{1}<t_{2}\cdots<t_{N}=T$, $n$ uniformly distributed over $\{1,2,\cdots,N-1\}$, $\lambda(\cdot)\in\mathbb{R}^{+}$ is a positive weighting function, and the expectation is taken with respect to $\bm{x}\sim p_{\rm{data}}$. $\bm{x}_{t_{n+1}}$ can be sampled using SDE (1) and $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}$ is calculated by $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}:=\Phi(\bm{x}_{t_{n+1}},t_{n+1},t_{n};\bm{\phi})$, where $\Phi(\cdots;\bm{\phi})$ represents the update function of a one-step ODE solver applied to the empirical PF ODE (3). Moreover, $\bm{\theta^{-}}$ is introduced to stabilize the training process and updated by an exponential moving average (EMA) strategy, i.e., given $10\leq\mu<$ 1, $\bm{\theta}^{-}\leftarrow\texttt{sg}(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta})$. With a well-trained model $\bm{\theta^{*}}$ such that $\mathcal{L}_{\rm{CD}}^{N}(\bm{\theta^{*}},\bm{\theta^{*}};\bm{\phi})=0$, and considering that the ODE solver $\Phi(\cdots;\bm{\bm{\phi}})$ has the local discretisation error, Theorem 1 in (Song et al., 2023) shows that the consistency distillation error is bounded:
$\mathop{\textnormal{sup}}\limits_{n,{\bm{x}}}\|{\bm{f}}_{\bm{\theta^{*}}}({\bm{x}},t_{n})-{\bm{f}}({\bm{x}},t_{n};\bm{\phi})\|_{2}=\mathcal{O}\left((\Delta t)^{p}\right),$ (9)
with $\Delta t$ and $p$ as defined in (Song et al., 2023).
@@
3.3 Latent Consistency Model
The Latent Consistency Model successfully extends consistency models (Song et al., 2023) to the image latent space corresponding with an auto-encoder (Rombach et al., 2022). To leverage the advantages of CMs, the authors proposed Latent Consistency Distillation, where the consistency function with condition information ${\bm{c}}$ can be defined as:
$\bm{f}(\bm{z}_{t},{\bm{c}},t)\mapsto\bm{z}_{0},$ (10)
which can directly predict the solution of PF-ODE for $t=0$. To leverage the knowledge of pretrained text-to-image models effectively, LCMs introduce the parameterisation method when parameterizing the consistency model with $\epsilon$-prediction teacher model:
$\tilde{{\bm{f}}}_{\bm{\theta}}({\bm{z}}_{t},{\bm{c}},t)=c_{\text{skip}}(t){\bm{z}}_{t}+c_{\text{out}}(t)\left(\frac{{\bm{z}}_{t}-\sigma_{t}\hat{\epsilon}_{\bm{\theta}}({\bm{z}}_{t},{\bm{c}},t)}{\alpha_{t}}\right),$ (11)
where the boundary condition satisfies $c_{\text{skip}}(0)=1$ and $c_{\text{out}}(0)=0$, $\alpha_{t}$ and $\sigma_{t}$ specify the noise schedule in the perturbation kernels. It’s worth noting that the above function, with appropriate parameterisation, is compatible with different prediction methods (e.g., $x$-prediction, $\epsilon$-prediction in (Ho et al., 2020), $v$-prediction in (Salimans & Ho, 2022)).
To accelerate distillation training, Luo et al. (2023a) proposed the skipping-step method to ensure self-consistency between timesteps $k$-steps apart, $t_{n+k}\rightarrow t_{n}$, instead of the original adjacent timesteps, $t_{n+1}\rightarrow t_{n}$. Additionally, Luo et al. (2023a) further implemented one-stage guided distillation (Meng et al., 2023) by solving the augmented PF-ODE. Then, the solution of $t_{n}$ from $t_{n+k}$ can be approximated with DDIM (Song et al., 2020a) or DPM-Solver (Lu et al., 2022a, b):
$\begin{split}\hat{{\bm{z}}}_{t_{n}}^{\phi,\omega}=(1+\omega)&\Phi({\bm{z}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})\\ &-\omega\Phi({\bm{z}}_{t_{n+k}},t_{n+k},t_{n},\varnothing;\bm{\phi}),\end{split}$ (12)
where $\Phi$ represents the update function of the PF-ODE solver with the pre-trained parameter $\phi$, and $\omega$ is the classifier-free guidance weight sampled from $\omega\in[\omega_{\text{min}},\omega_{\text{max}}]$.
The objective function of LCM is derived from the self-consistency property, similar to the consistency distillation loss in (Song et al., 2023):
$\mathcal{L}_{\rm{LCM}}(\bm{\theta},$ $\bm{\theta^{-}};\bm{\Phi}):=\mathbb{E}_{{\bm{z}},{\bm{c}},\omega,n}\bigg{[}$ (13)
$d\left(\tilde{\bm{f}}_{\bm{\theta}}({\bm{z}}_{t_{n+k}},{\bm{c}},t_{n+k});\tilde{\bm{f}}_{\bm{\theta}^{-}}(\hat{{\bm{z}}}_{t_{n}}^{\Phi,\omega},{\bm{c}},t_{n})\right)\bigg{]},$
where $d$ represents the distance between samples, with a common choice being the $\ell_{2}$ distance.